# CMB-PAInT: An inpainting tool for the cosmic microwave background

### v1

@@ 1 Introduction
One of the main issues to characterize the CMB anisotropies is how to deal with extragalactic and Galactic foreground emissions such as synchrotron, free-free, or thermal dust. These are astrophysical emissions between the last scattering surface and us. Precisely, the CMB has been measured in several frequency channels to exploit their different frequency response and to allow component-separation algorithms [3, 4, 5, 6, 7] to subtract an important part of them. Unfortunately, strongly contaminated regions such as the Galactic plane and the locations of extragalactic point sources can not be used for the statistical analyses, even after applying component-separation methods. In these regions, foregrounds are too large and complicated to be fully cleaned. The standard approach is to mask these regions using the confidence mask provided by each method, but this leads to loss of information and other technical difficulties. For instance, this prevents us from accurate characterization of large angular scales, which is crucial for the study of CMB anomalies [8, 9, 10, 11, 12, 13, 14, 15, 16]. In polarization the mask is even more problematic, not just at the harmonic level, but also at the map level. Masking $Q$ and $U$ Stokes parameters introduces an undesired mixing between $E$- and $B$-modes [17, 18], also referred as $E$-to-$B$ leakage. Actually, this is critical for the detection of the primordial $B$-mode signal as the $E$-mode signal is significantly larger. At map level, the lack of full-sky $Q$ and $U$ maps prevents us from having an accurate $E$- and $B$-mode reconstruction. The reason for this is the non-local nature of the transformations between $Q$ and $U$ and $E$ and $B$.
@@ 4.1 Constrained contribution
Figure 5: Dispersion of the inpainted maps estimated from the full set of 1200 inpainted realizations for the T (left panel), Q (middle) and U (right) components.
@@
In this section we study the 1-point probability distribution function (1-PDF) of the inpainted pixels. We also compute the 1-PDF for the $E$- and $B$-modes. Figures 7 and 8 provide the PDFs for the $T$, $Q$, and $U$ components and for the $E$- and $B$-modes, respectively, in the noiseless scenario. Similar results are obtained for the semi-realistic noise case, as shown in figure 9.
Figure 7: One-dimensional probability distribution inside the inpainted region for the $T$ (left panel), $Q$ (middle) and $U$ (right) components considering only the CMB signal. The black histogram corresponds to the input map. Green dots are the average value per bin obtained from the 1200 inpainted realizations. Orange and blue contours are the 68% and 95% C.I., respectively obtained for the distribution of the inpainted maps. Residuals are also shown in the lower panel.
@@ 4.3 $E$- & $B$-mode reconstruction
In order to assess the error in the $E$- and $B$-mode reconstructions, we compute the map of the standard deviation of the residuals. Starting from the 1200 $T$, $Q$ and $U$ inpainted realizations, we generate the corresponding $E$- and $B$-mode maps and compute the residuals by subtracting from them the input $E$- and $B$-mode. We calculate then the standard deviation maps, pixel by pixel, which are shown in Figure 11. On the one hand, for the $E$-mode, the maximum error outside the polarization common mask is 0.042 $\mu{K}$ which corresponds to around a 14 per cent of the typical amplitude of the E-mode signal, $\sigma_{E}\sim 0.29$ $\mu{K}$ at the considered resolution. On the other hand, for the $B$-mode, the maximum error is at the level of 0.019 $\mu{K}$, which is approximately the expected amplitude of the B-mode signal for r=0, $\sigma_{B}$. However, as figure 12 shows, the error decreases rapidly, and for 60% of the sky the maximum error is at the level of 20 per cent relative to $\sigma_{B}$. Concerning the $E$-mode, it is interesting to point out that the error is below 5% for more than the 71.7% of the sky, and then it goes down until it reaches a plateau, where we can not push the maximum residual to a lower value even if we extend the mask.
@@ 4.4 Power Spectra
Our final test is related to the power spectra estimation. Given the input $T$, $Q$ and $U$ full-sky maps, we calculate the $TT$, $EE$, $BB$ and $TE$ power spectrum, and compare it to the mean power spectra generated from the 1200 inpainted realizations. In particular, for this case where 30 per cent of the sky is inpainted, we find that the distribution of the values for each multipole of the inpainted realizations closely resembles a Gaussian distribution, even for $\ell<30$. For larger $f_{sky}$ to be inpainted, we expect the distribution to become more like a $\chi^{2}$. We also calculate the 68% and 95% C.L., as well as the residuals per multipole. The residuals are computed by taking the differences between input and median values over the 1200 realizations. Then, we divide by the upper or lower sigma [^7] Low sigma is computed by integrating from the median to the 16% of the low tail, i.e. it encapsulate 34% of the probability. The other 34% is in the upper sigma which is the integration between the median up to the 84%. For instance, if the residuals are positive, which means that the input value is below the median, we use the low sigma. to take into account possible asymmetries in the distributions, especially for low multipoles. Figure 13 shows the $TT$ and $TE$ power spectra, while in figure 14 $EE$ and $BB$ are plotted. Taking into account that we have a single CMB realization, the input spectra is noisy compared to the theoretical prediction due to the cosmic variance, both of them also plotted. Indeed, the mean power spectra obtained from the inpainted simulations follow that of the input CMB, rather than that of the theoretical model, finding that almost all recovered multipoles fall within the 95% C.L. Note that for the case of the B-mode, we are considering a scenario with r=0 (corresponding to the PR3 $\Lambda$CDM best fit model) and, therefore, only the contribution from lensing is present. However, we have checked that the power spectra is equally well recovered when starting with a simulation with $r$ different from zero.
@@
Figure 16 shows the results. As expected, for a deep transition in the edge of the mask, i.e. the case where pixels inside the mask are replaced by zeros, the $B$-mode power spectrum is completely dominated by the $E$-to-$B$ leakage (dark orange curve). This leakage can be reduced by one order of magnitude applying a diffuse inpainting as it smooths the discontinuity in the $Q$ and $U$ maps. However, the lensing signal is still hidden below the leakage. Our results show that the GCR is the best approach to recover the input power spectrum having residuals below the signal for all the multipole range. For comparison, we also consider the NaMaster pure $B$ approach, which recovers the B-mode power spectrum starting from the masked $Q$ and $U$ maps. For the considered case, this approach also fails reproducing the large angular scales of the B-mode.
@@ 5 CMB-PAInT
  • To compute the pixel covariance matrix either from an input angular power spectrum up to a certain $\ell_{max}$ or empirically from a set of simulations. Depending on the field to be inpainted, it can compute the covariance matrix of either $T$, $QU$, or $TQU$.
@@
This code can run on a NERSC-like cluster that uses slurm scheduling, or on a local machine, a Jupyter Notebook, or another cluster with different resource management. The only requirement is to consider the memory limitation. The code needs a configuration file that contains all the model and software parameters for running it. For example, the configuration file specifies whether to inpaint the polarization field or not, or whether to use an external .sh file. The code creates an instance of the CMB-PAInT class and runs one of the methods based on the user input: calculate_covariance, calculate_cholesky, or calculate_inpainting. If the methods run on a cluster without an input .sh file, it generates a .sh file based on NERSC with the resources from the configuration file, such as number of nodes, tasks, CPUs per task, time limits, email address, or partition. The code has two levels of parallelization. It distributes the work among a number of jobs ($N_{\mathrm{jobs}}$) that are submitted together, and each job uses mpi4py, the MPI standard for Python, to parallelize the assigned rows. This parallelization is used to calculate the covariance matrix from an input power spectrum and also to inpaint the maps. In the latter, the total number of realizations are distributed among the jobs and the tasks per job. Additionally, the code uses the dask package to perform Cholesky decomposition faster than the standard numpy implementation. The code saves intermediate products in a numpy file format, and Cholesky decomposition in a HDF5 binary data format.
@@ 6 Conclusions
In order to asses the performance of the methodology, a series of tests have been done in two different scenarios: (1) CMB signal only, and (2) CMB signal with semi-realistic noise simulations based on Planck Release 4. In particular, we checked that the constrained part of the inpainted maps was consistent with that expected from the model. We also studied the one-dimensional probability distribution of $T$, $Q$ and $U$, as well as those of the $E$- and $B$-modes. In both scenarios, they agreed well with the input values within the expected errors. Our methodology is also able to reconstruct accurately systematics and noise from an anisotropic field if they are included in the covariance matrix. However, this requires that all the correlations between pixels of the anisotropic field are perfectly characterised. Otherwise, artifacts would appear due to a mismatch between the statistical properties of the pixels outside the mask and the assumed model. This can be checked by inspecting the intermediate $z$ variables, which should follow a normal distribution if everything is consistent.
@@
Appendix A Appendix A: Convergence of the covariance matrix
@@
In section 4 we show the performance of the inpainting for two different scenarios: (1) CMB signal only and (2) CMB signal plus a semi-realistic noise realization. In the second case, the methodology is able to reconstruct accurately the systematics and noise. However, this is only possible if the full anisotropic covariance matrix is well characterized. In a more realistic case, where a limited number of simulations are used to compute the matrix numerically, the non-convergence of the matrix induces a mismatch between the matrix and the pixels outside the mask. This introduces some artifacts in the inpainted realizations. Figure 20 shows an example of the input and output $T$, $Q$ and $U$ maps using just 20 semi-realistic noise simulations to compute the covariance matrices. Even if the mismatch is mainly for the polarization field [^11] For temperature the noise is much smaller than the signal and, even if it were not well modelled in the covariance matrix, its effect is negligible as the error in the $z$ variables is masked by the regularization noise., strong cold and hot spots are induced in the inpainted temperature map through the TE correlation.
Figure 20: Example of a $T$ (left), $Q$ (middle) and $U$ (right) inpainted realization in the case where 20 semi-realistic noise simulations are used to characterize the noise covariance matrix. The top and bottom panels correspond to the input and inpainted maps, respectively.
Figure 21: Distribution of the $z$ variable associated to the $T$, $Q$ and $U$ fields for different number of simulations used to estimate the noise covariance matrix. For comparison, a Gaussian with the same standard deviation is also given in each case.
@@
The mismatch is even more clear in Figure 21, where the probability distribution of the $z$ variables (see eq. 2.12) is plotted for $T$, $Q$ and $U$. For comparison, a Gaussian curve with the same standard deviation is plotted. In all the cases shown in figure 21, the variable $z_{T}$, corresponding to the temperature field, follows a Gaussian distribution with zero mean and unit variance, $\mathcal{N}$(0, 1). This is because there is not a mismatch between the statistical properties of the pixels and those encoded in the matrix. In this case, even if we are not including correctly the noise and systematics on the matrix, the mismatch is simply masked by the regularization noise. However, the mismatch strongly affects the polarization field. In particular, since $Q$ and then $U$ are computed recursively, this effect is most notable for the $z$ variables associated to $U$, which is found to be the broadest one. The $z_{Q}$ and $z_{U}$ distributions become effectively $\mathcal{N}$(0, 1) when several thousand of noise simulations are considered to construct the matrix.

### v2

@@ 1 Introduction
One of the main issues to characterize the CMB anisotropies is how to deal with extragalactic and Galactic foreground emissions such as synchrotron, free-free, or thermal dust. These are astrophysical emissions between the last scattering surface and us. Precisely, the CMB has been measured in several frequency channels to exploit their different frequency response and to allow component-separation algorithms [3, 4, 5, 6, 7] to subtract an important part of them. Unfortunately, strongly contaminated regions such as the Galactic plane and the locations of extragalactic point sources can not be used for the statistical analyses, even after applying component-separation methods. In these regions, foregrounds are too large and complicated to be fully cleaned. The standard approach is to mask these regions using the confidence mask provided by each method, but this leads to loss of information and other technical difficulties. For instance, this prevents us from accurate characterization of large angular scales, which is crucial for the study of CMB anomalies [8, 9, 10, 11, 12, 13, 14, 15, 16]. In polarization the mask is even more problematic, not just at the harmonic level, but also at the map level. Masking $Q$ and $U$ Stokes parameters introduces an undesired mixing between $E$- and $B$-modes [17, 18], also referred as $E$-to-$B$ leakage. Actually, this is critical for the detection of the primordial $B$-mode signal as the $E$-mode signal is significantly larger. At map level, the lack of full-sky $Q$ and $U$ maps prevents us from having an accurate $E$- and $B$-mode reconstruction. The reason for this is the non-local nature of the transformations between $Q$ and $U$, and $E$ and $B$. Indeed, the main motivation of this work is an accurate reconstruction of the $E$- and $B$-mode outside the mask. These reconstructed maps are of interest for studies of isotropy and statistics, including the anomalies presented in the large angular scales. In order to characterise the quality of the inpainted maps, a few tests have been used including the power spectra. However, we note that for estimating the power spectra in masked skies other standard techniques are available which are described in the next paragraph.
@@ 4.1 Constrained contribution
Figure 5: Dispersion of the inpainted maps estimated from the full set of 1200 inpainted realizations for the T (left panel), Q (middle), and U (right) components.
@@
In this section we study the 1-point probability distribution function (1-PDF) of the inpainted pixels. We also compute the 1-PDF for the $E$- and $B$-modes. Figures 7 and 8 provide the PDFs for the $T$, $Q$, and $U$ components and the $E$- and $B$-modes, respectively, in the noiseless scenario. Similar results are obtained for the semi-realistic noise case, as shown in figure 9.
Figure 7: One-dimensional probability distribution inside the inpainted region for the $T$ (left panel), $Q$ (middle), and $U$ (right) components considering only the CMB signal. The black histogram corresponds to the input map. Green dots are the average value per bin obtained from the 1200 inpainted realizations. Orange and blue contours are the 68% and 95% C.L., respectively, obtained from the distribution of the inpainted maps. Residuals are also shown in the lower panel.
@@ 4.3 $E$- & $B$-mode reconstruction
In order to assess the error in the $E$- and $B$-mode reconstructions, we compute the map of the standard deviation of the residuals. Starting from the 1200 $T$, $Q$, and $U$ inpainted realizations, we generate the corresponding $E$- and $B$-mode maps and compute the residuals by subtracting from them the input $E$- and $B$-mode. We calculate then the standard deviation maps, pixel by pixel, which are shown in Figure 11. On the one hand, for the $E$-mode, the maximum error outside the polarization common mask is 0.042 $\mu{K}$ which corresponds to around a 14 per cent of the typical amplitude of the E-mode signal, $\sigma_{E}\sim 0.29$ $\mu{K}$ at the considered resolution. On the other hand, for the $B$-mode, the maximum error is at the level of 0.019 $\mu{K}$, which is approximately the expected amplitude of the B-mode signal for r=0, $\sigma_{B}$. However, as figure 12 shows, the error decreases rapidly, and for 60% of the sky the maximum error is at the level of 20 per cent relative to $\sigma_{B}$. Concerning the $E$-mode, it is interesting to point out that the error is below 5% for more than the 71.7% of the sky, and then it goes down until it reaches a plateau, where we can not push the maximum residual to a lower value even if we extend the mask.
@@ 4.4 Power Spectra
Our final test is related to the power spectra estimation. Given the input $T$, $Q$, and $U$ full-sky maps, we calculate the $TT$, $EE$, $BB$ and $TE$ power spectrum, and compare it to the mean power spectra generated from the 1200 inpainted realizations. In particular, for this case where 30 per cent of the sky is inpainted, we find that the distribution of the values for each multipole of the inpainted realizations closely resembles a Gaussian distribution, even for $\ell<30$. For larger $f_{sky}$ to be inpainted, we expect the distribution to become more like a $\chi^{2}$. We also calculate the 68% and 95% C.L., as well as the residuals per multipole. The residuals are computed by taking the differences between input and median values over the 1200 realizations. Then, we divide by the upper or lower sigma [^7] Low sigma is computed by integrating from the median to the 16% of the low tail, i.e. it encapsulate 34% of the probability. The other 34% is in the upper sigma which is the integration between the median up to the 84%. For instance, if the residuals are positive, which means that the input value is below the median, we use the low sigma. to take into account possible asymmetries in the distributions, especially for low multipoles. Figure 13 shows the $TT$ and $TE$ power spectra, while in figure 14 $EE$ and $BB$ are plotted. Taking into account that we have a single CMB realization, the input spectra is noisy compared to the theoretical prediction due to the cosmic variance, both of them also plotted. Indeed, the mean power spectra obtained from the inpainted simulations follow that of the input CMB, rather than that of the theoretical model, finding that almost all recovered multipoles fall within the 95% C.L. Note that for the case of the B-mode, we are considering a scenario with r=0 (corresponding to the PR3 $\Lambda$CDM best fit model) and, therefore, only the contribution from lensing is present. However, we have checked that the power spectra is equally well recovered when starting with a simulation with $r$ different from zero.
@@
Figure 16 shows the results. As expected, for a deep transition in the edge of the mask, i.e. the case where pixels inside the mask are replaced by zeros, the $B$-mode power spectrum is completely dominated by the $E$-to-$B$ leakage (dark orange curve). This leakage can be reduced by one order of magnitude applying a diffuse inpainting as it smooths the discontinuity in the $Q$ and $U$ maps. However, the lensing signal is still hidden below the leakage. Our results show that the GCR is the best approach to recover the input power spectrum having residuals below the signal for all the multipole range. For comparison, we also consider the NaMaster pure $B$ approach, which recovers the B-mode power spectrum starting from the masked $Q$ and $U$ maps. For the considered case, this approach also fails reproducing the large angular scales of the B-mode. These results show that obtaining the power spectra from an inpainted map could be used as an alternative to more standard methods. However, further work is needed to validate the usefulness of this approach.
@@ 5 CMB-PAInT
  • To compute the pixel covariance matrix from an input angular power spectrum up to a certain $\ell_{max}$. Depending on the field to be inpainted, it can compute the covariance matrix of either $T$, $QU$, or $TQU$.
@@
This code can run on a NERSC-like cluster that uses slurm scheduling, or on a local machine, a Jupyter Notebook, or another cluster with different resource management. The only requirement is to consider the memory limitation. The code needs a configuration file that contains all the model and software parameters for running it. For example, the configuration file specifies whether to inpaint the polarization field or not, or whether to use an external .sh file. The code creates an instance of the CMB-PAInT class and runs one of the methods based on the user input: calculate_covariance, calculate_cholesky, or calculate_inpainting. If the methods run on a cluster without an input .sh file, it generates a .sh file based on NERSC with the resources from the configuration file, such as number of nodes, tasks, CPUs per task, time limits, email address, or partition. The code has two levels of parallelization. It distributes the work among a number of jobs ($N_{\mathrm{jobs}}$) that are submitted together, and each job uses mpi4py, the MPI standard for Python, to parallelize the assigned rows or maps. This parallelization is used to calculate the covariance matrix from an input power spectrum and also to inpaint the maps. In the latter, the total number of realizations are distributed among the jobs and the tasks per job. Additionally, the code uses the dask package to perform Cholesky decomposition faster than the standard numpy implementation. The code saves intermediate products in a numpy file format, and Cholesky decomposition in a HDF5 binary data format.
@@ 6 Conclusions
In order to asses the performance of the methodology, a series of tests have been done in two different scenarios: (1) CMB signal only, and (2) CMB signal with semi-realistic noise simulations based on Planck Release 4. In particular, we checked that the constrained part of the inpainted maps was consistent with that expected from the model. We also studied the one-dimensional probability distribution of $T$, $Q$, and $U$, as well as those of the $E$- and $B$-modes. In both scenarios, they agreed well with the input values within the expected errors. Our methodology is also able to reconstruct accurately systematics and noise from an anisotropic field if they are included in the covariance matrix. However, this requires that all the correlations between pixels of the anisotropic field are perfectly characterised. Otherwise, artifacts would appear due to a mismatch between the statistical properties of the pixels outside the mask and the assumed model. This can be checked by inspecting the intermediate $z$ variables, which should follow a normal distribution if everything is consistent.
@@
Appendix A Convergence of the covariance matrix
@@
In section 4 we show the performance of the inpainting for two different scenarios: (1) CMB signal only and (2) CMB signal plus a semi-realistic noise realization. In the second case, the methodology is able to reconstruct accurately the systematics and noise. However, this is only possible if the full anisotropic covariance matrix is well characterized. In a more realistic case, where a limited number of simulations are used to compute the matrix numerically, the non-convergence of the matrix induces a mismatch between the matrix and the pixels outside the mask. This introduces some artifacts in the inpainted realizations. Figure 20 shows an example of the input and output $T$, $Q$, and $U$ maps using just 20 semi-realistic noise simulations to compute the covariance matrices. Even if the mismatch is mainly for the polarization field [^11] For temperature the noise is much smaller than the signal and, even if it were not well modelled in the covariance matrix, its effect is negligible as the error in the $z$ variables is masked by the regularization noise., strong cold and hot spots are induced in the inpainted temperature map through the TE correlation.
@@
The mismatch is even more clear in Figure 21, where the probability distribution of the $z$ variables (see eq. 2.12) is plotted for $T$, $Q$, and $U$. For comparison, a Gaussian curve with the same standard deviation is plotted. In all the cases shown in figure 21, the variable $z_{T}$, corresponding to the temperature field, follows a Gaussian distribution with zero mean and unit variance, $\mathcal{N}$(0, 1). This is because there is not a mismatch between the statistical properties of the pixels and those encoded in the matrix. In this case, even if we are not including correctly the noise and systematics on the matrix, the mismatch is simply masked by the regularization noise. However, the mismatch strongly affects the polarization field. In particular, since $Q$ and then $U$ are computed recursively, this effect is most notable for the $z$ variables associated to $U$, which is found to be the broadest one. The $z_{Q}$ and $z_{U}$ distributions become effectively $\mathcal{N}$(0, 1) when several thousand of noise simulations are considered to construct the matrix.
Figure 20: Example of a $T$ (left), $Q$ (middle), and $U$ (right) inpainted realization in the case where 20 semi-realistic noise simulations are used to characterize the noise covariance matrix. The top and bottom panels correspond to the input and inpainted maps, respectively.
Figure 21: Distribution of the $z$ variable associated to the $T$, $Q$, and $U$ fields for different number of simulations used to estimate the noise covariance matrix. For comparison, a Gaussian with the same standard deviation is also given in each case.
Appendix B Robustness against the model
In section 4 we generate a pixel covariance matrix that match perfectly the input simulation. In a real situation this is not possible. Here, we introduce a small mismatch between the model used for the input simulation and the one used for the pixel covariance matrix estimation. As described in 2, we use the best fit to the $\Lambda$CDM model to generate the input simulation (only CMB is included in the simulation). We introduce a small deviation in the input parameters taking into account the correlation matrix of the estimated parameters inferred from the Planck 2018 data. In Table 1 we show the best fit parameters and the modified ones. The alternative parameters were obtained as a Gaussian random realisation of the parameters centred in the best-fit model and following the correlation matrix. Therefore, they are also consistent with the Planck data within the estimated errors.
$\Lambda CDM$ best fit Modified $\Lambda CDM$
$\Omega_{b}h^{2}$ 0.02238280 0.02244959
$\Omega_{c}h^{2}$ 0.1201075 0.1194516
$H_{0}$ 67.32117 67.57953
$\tau$ 0.05430842 0.05744905
${\rm{ln}}(10^{10}A_{s})$ 3.044784 3.056551
$n_{s}$ 0.9660499 0.9654136
Table 1: Cosmological parameters. Left: $\Lambda CDM$ best fit. Model used for the input simulation. Right: Modified model according to the Planck 2018 errorbars and correlations.
We compute the relative error with respect cosmic variance in the $E$- and $B$-mode reconstruction using the 1200 inpainted realizations, as explained in section 4.3, for both cases, the exact and the modified model. Then, we get the differences between previous errors, which are on average at the level of 0.1% (errors are slightly larger in the modified case) as it is shown in Figure 22.
Figure 22: Distribution of differences of the errors relative to the cosmic variance of the exact minus the modified model for $E$-mode (left) and $B$-mode (right).
[…]
Similar results are found at the power spectra level. This is expected taking into account that the differences between models in this multipole range ($\ell$ = 2-192) is much smaller than the cosmic variance, so smaller than the differences between different realizations of the same model.
As an extra test, we consider a special scenario where tensor-to-scalar ratio (r) is equal to 0.004. In this case, most of the constraining power is encoded in the $B$-mode large scales, so differences in the model can impact and bias the r estimation. We demonstrate that an iterative process can be a good approach to this issue. Taking into account that tensorial modes have not been detected yet, it is reasonable to start with a model where r = 0. We generate 1200 inpainted realizations assuming r = 0 model and we estimate for each of them the posterior (using a flat prior) of r using the following exact likelihood in harmonic space:
[…]
We finally show a comparison between the $B$-mode power spectrum obtained using different models. Left panel of Figure 24 shows $D_{\ell}^{BB}$ of the input simulation (red), the theoretical curve of the input model (black), and the estimated $\tilde{D}_{\ell}^{BB}$ for the different fiducial models (corresponding to different iterations), which are described by the median of the distribution of the 1200 spectra, obtained from each of the inpainted realizations, and the 68% C.L (error bars). In the right panel we show $\beta_{\ell}$ defined as,
$\beta_{\ell}=\frac{x_{\ell}}{\sigma_{\ell}}$ (B.2)
where $x_{\ell}$ is the difference between the input spectrum and the median, and $\sigma_{\ell}$ is the 68% two sided C.L, thus allowing asymmetric error bars for low $\ell$. It is clear that a bad choice of the model has a bigger impact on the first multipoles, corresponding to the reionization bump, than in the lensing dominated range.
Figure 24: Left panel: $D_{\ell}^{BB}$ for the input simulation (in red), theoretical curve for r = 0.004 (in black), and $\tilde{D}_{\ell}^{BB}$ for different fiducial models, obtained from the distribution of 1200 inpainted realizations. Error bars are also derived from the 68% C.L of that distribution. Right panel: $\beta_{\ell}$ for different models.